# Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach

### v1

@@ 6 Application to Microcredit
In 2022, microcredit reached more than 170 million borrowers worldwide, with a total loan portfolio of over \$140 billion (Convergences, 2023). Theory suggests that microborrowing can be beneficial or harmful (Banerjee, 2013, Garz et al., 2021), and evidence suggests that average effects are small if not negligible (Banerjee et al., 2015b, Meager, 2019). Analysis of heterogeneity mostly relied on rank preservation, and suggested mainly positive impact in the upper tail of income/profits distributions (e.g., Meager, 2022). Yet, evidence of negative effects has been scarce. I revisit five studies which randomized offering a microloan at the individual level [^4] The only exception is Egypt, in which the treated group was offered a loan size twice as large as the offered in the control group., and find evidence of both positive and negative treatment effects of microcredit. A brief description of the datasets is presented in Table 2.
CAIDE is able to provide informative bounds, even in this challenging context. First because the distributions of income/profits are similar between treated and control groups. This is reflected in not statistically significant ATEs in Table 2, and estimated $[\theta_{L},\theta_{U}]$ close to $[0,1]$ if covariates are not considered (Table 3). Second, because the sample sizes are relatively small (ranging from 548 to 1964), with a large number of covariates (from 24 to 97). Finally, because predicting business outcomes is itself a difficult task (McKenzie and Sansone, 2019).
@@
Using $K=10$, I fit the estimators of Section 4555The sample-splitting estimators of Section 3 with a 50-50 split lead all confidence intervals to be $[0,1]$ at $10\%$ significance level.. For each fold, I fit another layer of sample splitting to pick via $10-$fold cross-validation one of the models: Quantile Forests, Quantile Neural Networks, Random Forest (Regression), Neural Networks (Reg.), Elastic Net (Reg.), Extreme Gradient Boosting (Reg.), and Constant ($s(X)=0)$. Technical details are provided in Appendix C.
@@
Table 3 shows the estimates for the lower and upper bounds on $\theta$ for $\delta=0$ using CAIDE with and without covariates [^6] Note that the estimators and one-sided CIs without covariates ($s(x)=0$) are equivalent to the ones in Fan and Park (2010).. The outcome is the logarithm of annual household income. Considering that the outcome is continuous, it follows that $\theta=P(Y(1)-Y(0)\leq 0)=P(Y(1)-Y(0)<0)$, the lower bound on $\theta$ is the minimum proportion of households that have a negative effect from the microloan, and the upper bound is the maximum proportion. Note that one minus the upper bound denotes the minimum proportion of households that have a positive effect. Results for $\delta=-0.05$ and $\delta=0.05$ are similar and are presented in Table 4. In Bosnia, for example, I estimate that at least $10.8\%$ of the households would be better off without the microloan (p-value $<10^{-3}$), and at most $94.6\%$ have negative effects (p-value $0.028$). That is, at least $6.4\%$ are better off ($1-0.946$). Overall, I find significant lower bounds at the $10\%$ level for all of the five datasets, and at the $5\%$ level for four of them. As for the upper bound, I find statistical significance at $5\%$ in four of the datasets. These results reveal important presence of heterogeneity, even when the ATEs are not statistically significant (Table 2).
@@
Incorporating covariates makes substantial difference in many of the point estimates. In Egypt, for example, the lower bound without covariates is $0.009$ (p-value $0.264$), whereas incoporating covariates improves the estimate to $0.260$ (p-value $<10^{-3}$). The differences between the point estimates with and without covariates are siginificant at the $5\%$ level for both lower and upper bounds for Egypt, Philippines (2) and All.
@@
Table 3 also illustrates how CAIDE can outperform SJLS (Semenova, 2023, Ji et al., 2023) in terms of narrower confidence intervals. In the case of the lower bound for Bosnia, for example, it is estimated $10.8\%$ with CAIDE vs $2.4\%$ with SJLS, leading to substancial difference in p-values ($<10^{-3}$ vs $0.254$). It also illustrates how SJLS can estimate negative lower bounds (South Africa), or upper bounds that are greater than one (Philippines (1)). In both cases, the estimates with CAIDE are statistically significant at the $10\%$ level.
@@
Table 5 in Online Appendix IV shows results for business profits. Again, including covariates improves the bounds. Averaging estimates from all datasets, I find that at least $10.8\%$ (p-value $<10^{-3}$) of the population was harmed, and at least $9.8\%$ (p-value $<10^{-3}$) benefited from microcredit.
@@
An important note is that this evidence does not inform the nature of the negative effects or their magnitude. It could be the case, for example, that the negative outcomes are relatively small and are a natural consequence of investments being risky. In that case, the decision to take a microloan could still be rational if the prospect of gains is better than the risk of loss.
@@ 7 Conclusion
I propose two novel inference approaches to the distributional treatment effect $\theta=P(Y(1)-Y(0)\leq\delta)$, for any fixed $\delta$. A new characterization of the sharp identified set of $\theta$ is provided, which is based on covariate adjustment. It has the desirable property of being robust to the choice of the covariate-adjustment term $s(x)$, in the sense that if this term is misspecified, the identified set is wider but still valid. As a consequence, any method can be used to estimate $s$, including machine learning algorithms. I provide finite-sample valid inference using sample-splitting and asymptotic inference using cross-fitting, which is also asymptotically exact under additional conditions.
@@
The practical relevance of the method is illustrated in a simulation study and in an application to microcredit. The application reveals important presence of heterogeneity, with significant negative and positive treatment effects. The results show that the method is able to provide informative bounds even in challenging contexts, such as when the ATE is not statistically significant, the sample size is small, and the number of covariates is large.
@@
By using CAIDE, an applied researcher can learn more about heterogeneous effects not only by gathering larger samples, but also by collecting information on covariates that help predict potential outcomes.
Appendix A Comparison with Semenova (2023) and Ji et al. (2023)
@@
I compare the asymptotic properties of CAIDE with the estimators proposed by Semenova (2023) and Ji et al. (2023) (denoted by SJLS for the first letter of the names of the authors). I focus on the cross-fitting version of the estimators (Definition 2(iii)), and compare them to the cross-fitting estimators in SJLS. Note that in the case of the DTE $\theta$, the estimators of Semenova (2023) and Ji et al. (2023) are equivalent.
@@
Since SJLS consider a known propensity score, I compare SJLS to a version of CAIDE that uses the true propensity score (Online Appendix I) instead of using the fraction of treated units in the sample. In Appendix B I show that using the actual number of treated units (as in Definition 2(iii)) instead of the known propensity score leads to the same asymptotic distribution but with a smaller asymptotic variance. Hence, using the true propensity score in CAIDE is the natural comparison. I focus on the lower bound for simplicity, and analogous results hold for the upper bound. Define
$\widehat{\Delta}(t)=\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{D_{i}}{p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)-\frac{1-D_{i}}{1-p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\biggr{)}$
@@
Then, CAIDE is equivalent to $\hat{\theta}_{C}:=\max_{t}\widehat{\Delta}(t)$ and SJLS is given by $\hat{\theta}_{S}:=\widehat{\Delta}(0)$. Theorem 3 of this paper is comparable to Theorem 4.1 in Semenova (2023), and Theorem 3.1 and Proposition 3.2 in Ji et al. (2023). They all show asymptotic validity of confidence intervals. Define confidence intervals as $CI_{C}$ for CAIDE and $CI_{S}$ for SJLS. Since by construction $\hat{\theta}_{C}\geq\hat{\theta}_{S}$, it follows that $\hat{\theta}_{C}$ first order stochastically dominates $\hat{\theta}_{S}$. Under the conditions of these theorem, it follows that, for any sequence $\theta_{n}<\theta_{P_{n}}^{*}$, where $\theta_{P_{n}}^{*}$ is the sharp lower bound as in (2.5),
$\liminf_{n\to\infty}P_{n}(\theta_{n}\in CI_{S})-P(\theta_{n}\in CI_{C})\geq 0.$
Hence, $CI_{C}$ has more power than $CI_{S}$ in the sense that $CI_{S}$ may include elements below the sharp lower bound that are not included in $CI_{C}$. Note that validity of $CI_{C}$ relies on A2(iii), on the uniqueness of the maximizer of $\widehat{\Delta}(t)$, which is not assumed in SJLS.
Appendix B Estimated vs True Propensity Score
@@
I show that, when the propensity score is constant, the asymptotic variances of the estimators in 2(iii) are smaller than those of alternative estimators that consider the true propensity score. When the propensity score is known and not constant as in A1.ii, an alternative estimator that uses the true propensity score is proposed in Online Appendix I-B. The lower bound, for example, is given by
@@
Since the third term above is never negative and in general different from zero, the asymptotic variance of $\widehat{\theta}_{A}$ is smaller than the one of $\widehat{\theta}^{B}_{A}$.
@@
Appendix C Estimation of the covariate-adjustment term
I consider two approaches to estimate the covariate-adjustment term $s(x)$. I focus the explanation on the case of the lower bound for simplicity. First, I estimate $\hat{F}_{j}(t|x)$ ($j=0,1$) with one of two methods, and then I calculate $\hat{s}$ from $\hat{F}_{j}(t|x)$.
The first method for estimating $\hat{F}_{j}(t|x)$ is to use quantile models, such as quantile forests or quantile neural networks. They give estimates for quantiles, that is, $\hat{F}^{-1}_{j}(\tau|x)$, for all values of $x$ in the sample. $\hat{F}_{j}(t|x)$ can be calculated from $\hat{F}^{-1}_{j}(\tau|x)$ by linear interpolation. In Sections 5 and 6, I consider a grid for $\tau$ from 00 to $1$ by increments of $0.01$, i.e. $\tau_{k}\in\{0,0.01,\dots,1\}$. For an arbitrary $t$, pick $\tau_{a}$ and $\tau_{b}$ such that $t_{a}=\hat{F}^{-1}_{j}(\tau_{a}|x),t_{b}=\hat{F}^{-1}_{j}(\tau_{b}|x)$ are the two closest values to $t$ such that $t_{a}\leq t\leq t_{b}$. Then, $\hat{F}_{j}(t|x)=\tau_{a}+\frac{\tau_{b}-\tau_{a}}{t_{b}-t_{a}}(t-t_{a})$.
@@ Appendix C Estimation of the covariate-adjustment term
I fit all models in R. For quantile neural networks, I use the package qrnn with 3 hidden nodes, in 1 trial and maximum of 100 iterations. Other parameters are the default choice. For quantile random forests, I use the package ranger with 1,000 trees and otherwise default hyperparameters. For all the regression methods, I use the package mlr3 with default values of hyperparameters. The packages for all models, used through mlr3 are: random forests (ranger), extreme gradient boosting (xgboost), support vector machine (e1071), neural networks (nnet), elastic net (glmnet). The constant/none model simply takes $s(x)=0$.
For choosing the best model in the application of Section 6, I use cross validation inside each fold. That is, for each fold, I fit all the models in a 10-fold cross-validation, and choose the model that gives the largest (smallest) lower (upper) bound. Note that the test set is never used in this approach.
@@
Appendix D Definitions Ommited from the Main Text
@@
Appendix E Proofs of the Main Results
@@
$\hskip 20.0pt-P\left(\hat{F}_{0,L}(\hat{t}_{max})-F_{0,L}(\hat{t}_{max})>c_{\alpha,0}\right)$ (E.1)
@@
$\hskip 15.0pt-\mathbb{E}\left[P\left(\sup_{t}\left(\hat{F}_{0,L}(t)-F_{0,L}(t)\right)>c_{\alpha,0}\Bigm{|}\mathcal{I}_{A}\right)\right]$ (E.2)
$\geq 1-\mathbb{E}\left[\exp{(-2|\mathcal{I}^{M}_{1}|c_{\alpha,1}^{2})}\right]-\mathbb{E}\left[\exp{(-2|\mathcal{I}^{M}_{0}|c_{\alpha,0}^{2})}\right]$ (E.3)
@@
(E.1) holds since $\hat{F}_{0,L}(t)-F_{0,L}(t)$ is symmetric around zero for all $t$. (E.2) holds by the law of iterated expectations, and note the expected values are in respect to $\mathcal{I}_{A}$. (E.3) holds by the one-sided DKW inequality as in Massart (1990). Note the use of sample splitting and conditioning on the auxiliary set $\mathcal{I}_{A}$ is important since it ensures that $(Y_{i},\hat{s}_{L}(X_{i}),D_{i})_{i\in\mathcal{I}^{1}_{M}\cup\mathcal{I}^{0}_{M}}$ are iid, assumption required for the DKW inequality.
@@
Theorem 3 is restated as Theorem E.1:
Theorem E.1 (Asymptotic distribution of cross-fitting estimators).
@@
$=\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(\hat{t}_{A},s_{A,P_{n}})-F_{P_{n},j}(\hat{t}_{A},s_{A,P_{n}})\right)+o_{P_{n}}(1)$ (E.4)
$=\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(t_{A,P_{n}},s_{A,P_{n}})-F_{P_{n},j}(t_{A,P_{n}},s_{A,P_{n}})\right)+o_{P_{n}}(1)$ (E.5)

### v2

@@ 6 Application to Microcredit
In 2022, microcredit reached more than 170 million borrowers worldwide, with a total loan portfolio of over \$140 billion (Convergences, 2023). Theory suggests that microborrowing can be beneficial or harmful (Banerjee, 2013, Garz et al., 2021), and evidence suggests that average effects are small if not negligible (Banerjee et al., 2015b, Meager, 2019). Analysis of heterogeneity mostly relied on rank preservation and suggested mainly positive impact in the upper tail of income/profits distributions (e.g., Meager, 2022). Yet, evidence of negative effects has been scarce. I revisit five studies that randomized offering a microloan at the individual level [^4] The only exception is Egypt, in which the treated group was offered a loan size twice as large as the one offered in the control group., and find evidence of both positive and negative treatment effects of microcredit. Table 2 presents a brief description of the datasets.
CAIDE is able to provide informative bounds, even in this challenging context. First, the setting is challenging because the distributions of income/profits are similar between the treated and control groups. This similarity is reflected in not statistically significant ATEs in Table 2, and estimated $[\theta_{L},\theta_{U}]$ close to $[0,1]$ if covariates are not considered (Table 3). Second, because the sample sizes are relatively small (ranging from 548 to 1964), with a large number of covariates (from 24 to 97). Finally, because predicting business outcomes is itself a difficult task (McKenzie and Sansone, 2019).
@@
Using $K=10$, I fit the estimators of Section 4555The sample-splitting estimators of Section 3 with a 50-50 split lead all confidence intervals to be $[0,1]$ at $10\%$ significance level.. For each fold, I fit another layer of sample splitting to pick via $10-$fold cross-validation one of the models: Quantile Forests, Quantile Neural Networks, Random Forest (Regression), Neural Networks (Reg.), Elastic Net (Reg.), Extreme Gradient Boosting (Reg.), and Constant ($s(X)=0)$. Technical details are provided in Appendix A.3.
@@
Table 3 shows the estimates for the lower and upper bounds on $\theta$ for $\delta=0$ using CAIDE with and without covariates [^6] Note that the estimators and one-sided CIs without covariates ($s(x)=0$) are equivalent to the ones in Fan and Park (2010).. The outcome is the logarithm of annual household income. Considering that the outcome is continuous, it follows that $\theta=P(Y(1)-Y(0)\leq 0)=P(Y(1)-Y(0)<0)$, the lower bound on $\theta$ is the minimum proportion of households that have a negative effect from the microloan, and the upper bound is the maximum proportion. Note that one minus the upper bound denotes the minimum proportion of households with a positive effect. Results for $\delta=-0.05$ and $\delta=0.05$ are similar and are presented in Table 4. In Bosnia, for example, I estimate that at least $10.8\%$ of the households would be better off without the microloan (p-value $<10^{-3}$), and at most $94.6\%$ have negative effects (p-value $0.028$). That is, at least $6.4\%$ are better off ($1-0.946$). Overall, I find significant lower bounds at the $10\%$ level for all five datasets and at the $5\%$ level for four of them. As for the upper bound, I find statistical significance at $5\%$ in four datasets. These results reveal an important presence of heterogeneity, even when the ATEs are not statistically significant (Table 2).
@@
Incorporating covariates makes a substantial difference in many of the point estimates. In Egypt, for example, the lower bound without covariates is $0.009$ (p-value $0.264$), whereas incorporating covariates improves the estimate to $0.260$ (p-value $<10^{-3}$). The differences between the point estimates with and without covariates are significant at the $5\%$ level for both the lower and upper bounds for Egypt, the Philippines (2) and All.
@@
Table 3 also illustrates how CAIDE can outperform SJLS (Semenova, 2023, Ji et al., 2023) in terms of narrower confidence intervals. In the case of the lower bound for Bosnia, for example, it is estimated $10.8\%$ with CAIDE vs. $2.4\%$ with SJLS, leading to a substantial difference in p-values ($<10^{-3}$ vs. $0.254$). It also illustrates how SJLS can estimate negative lower bounds (South Africa), or upper bounds greater than one (Philippines (1)). In both cases, the estimates with CAIDE are statistically significant at the $10\%$ level.
@@
Table 5 in Online Appendix OA.4 shows results for business profits. Again, including covariates improves the bounds. Averaging estimates from all datasets, I find that at least $10.8\%$ (p-value $<10^{-3}$) of the population was harmed, and at least $9.8\%$ (p-value $<10^{-3}$) benefited from microcredit.
@@
An important note is that this evidence does not inform the nature of the negative effects or their magnitude. For example, the negative outcomes could be relatively small and a natural consequence of investments being risky. In that case, the decision to take a microloan could still be rational if the prospect of gains is better than the risk of loss.
@@ 7 Conclusion
I propose two novel inference approaches to the distributional treatment effect $\theta=P(Y(1)-Y(0)\leq\delta)$ for any fixed $\delta$. A new characterization of the sharp identified set of $\theta$ is provided, based on covariate adjustment. It has the desirable property of being robust to the choice of the covariate-adjustment term $s(x)$. If this term is misspecified, the identified set is wider but still valid. As a consequence, any method can be used to estimate $s$, including machine learning algorithms. I provide finite-sample valid inference using sample-splitting and asymptotic inference using cross-fitting, which is also asymptotically exact under additional conditions.
@@
I illustrate the practical relevance of the method in a simulation study and an application to microcredit. The application reveals an important presence of heterogeneity, with significant negative and positive treatment effects. The results show that the method is able to provide informative bounds even in challenging contexts, such as when the ATE is not statistically significant, the sample size is small, and the number of covariates is large.
@@
By using CAIDE, an applied researcher can learn more about heterogeneous effects not only by gathering larger samples but also by collecting information on covariates that help predict potential outcomes.
Appendix A
A.1 Comparison with Semenova (2023) and Ji et al. (2023)
@@
I compare the asymptotic properties of CAIDE with the estimators proposed by Semenova (2023) and Ji et al. (2023) (denoted by SJLS for the first letter of the names of the authors). I focus on the cross-fitting version of my estimators (Definition 2(iii)) and compare them to the cross-fitting estimators in SJLS. Note that in the case of the DTE, $\theta$, the estimators of Semenova (2023) and Ji et al. (2023) are equivalent.
@@
Since SJLS considered a known propensity score, I compare SJLS to a version of CAIDE that uses the true propensity score (Online Appendix OA.1) instead of using the fraction of treated units in the sample. In Appendix A.2, I show that using the actual number of treated units (as in Definition 2(iii)) instead of the known propensity score leads to a smaller asymptotic variance. Hence, using the true propensity score in CAIDE is the natural comparison. I focus on the lower bound for simplicity, and analogous results hold for the upper bound. Define
$\widehat{\Delta}(t,\hat{s}_{L})=\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{D_{i}}{p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)-\frac{1-D_{i}}{1-p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\biggr{)},$
where $\hat{s}_{L}=(\hat{s}_{L,1},\dots,\hat{s}_{L,K})$.
@@
Then, CAIDE is equivalent to $\hat{\theta}_{C}:=\max_{t}\widehat{\Delta}(t,\hat{s}_{L})$ and SJLS is given by $\hat{\theta}_{S}:=\widehat{\Delta}(0,\hat{s}_{L})$. Theorem 3 of this paper is comparable to Theorem 4.1 in Semenova (2023), and Theorem 3.1 and Proposition 3.2 in Ji et al. (2023). They all show asymptotic validity of confidence intervals. Define confidence intervals as $CI_{C}$ for CAIDE and $CI_{S}$ for SJLS. Since by construction $\hat{\theta}_{C}\geq\hat{\theta}_{S}$, it follows that $\hat{\theta}_{C}$ first order stochastically dominates $\hat{\theta}_{S}$. Let $\theta_{n}^{*}=\Delta_{P_{n}}(0,s^{*}_{L})$ be the sharp lower bound as in (2.5) and Theorem 1. Under the conditions of these theorems, it follows that, for any sequence $\theta_{n}<\theta_{n}^{*}$,
$\liminf_{n\to\infty}P_{n}(\theta_{n}\in CI_{S})-P_{n}(\theta_{n}\in CI_{C})\geq 0.$
Furthermore, consider a scenario where $\max_{t}\Delta_{P_{n}}(t,s_{L,P})>\Delta_{P_{n}}(0,s_{L,P})+M$ for some $M>0$, where $s_{L,P}$ is the limit of $\hat{s}_{L,k}$ (as in A2(ii)). This is usually the case when $s_{L,P}$ is not a sharp transformation $s^{*}_{L,P}$. Then, for any $\Delta_{P_{n}}(0,s_{L,P_{n}})+M\leq\theta_{n}\leq\max_{t}\Delta_{P_{n}}(t,s_{L,P_{n}})$, the inequality above holds with strict inequality.
Since $\theta_{n}<\theta_{n}^{*}$, a lower coverage probability translates into higher power. Hence, $CI_{C}$ dominates $CI_{S}$ in terms of power. Note that validity of $CI_{C}$ relies on A2(iii), on the uniqueness of the maximizer of $\Delta_{P_{n}}(t,s_{L,P})$, which is not assumed in SJLS. I discuss two approaches to weaken this assumption in Online Appendix OA.3.
A.2 Estimated vs True Propensity Score
@@
I show that when the propensity score is constant, the asymptotic variances of the estimators in 2(iii) are smaller than those of alternative estimators that consider the true propensity score. When the propensity score is known and not constant as in A1.ii, an alternative estimator that uses the true propensity score is proposed in Online Appendix OA.1-B. The lower bound, for example, is given by
@@
Since the third term above is never negative and, in general, different from zero, the asymptotic variance of $\widehat{\theta}_{A}$ is smaller than the one of $\widehat{\theta}^{B}_{A}$.
@@
A.3 Estimation of the covariate-adjustment term
I consider two approaches to estimate the covariate-adjustment term $s(x)$. For simplicity, I focus the explanation on the case of the lower bound. First, I estimate $\hat{F}_{j}(t|x)$ ($j=0,1$) with one of two methods, and then I calculate $\hat{s}$ from $\hat{F}_{j}(t|x)$.
The first method for estimating $\hat{F}_{j}(t|x)$ is using quantile models, such as quantile forests or neural networks. They give estimates for quantiles, that is, $\hat{F}^{-1}_{j}(\tau|x)$, for all values of $x$ in the sample. $\hat{F}_{j}(t|x)$ can be calculated from $\hat{F}^{-1}_{j}(\tau|x)$ by linear interpolation. In Sections 5 and 6, I consider a grid for $\tau$ from 00 to $1$ by increments of $0.01$, i.e. $\tau_{k}\in\{0,0.01,\dots,1\}$. For an arbitrary $t$, pick $\tau_{a}$ and $\tau_{b}$ such that $t_{a}=\hat{F}^{-1}_{j}(\tau_{a}|x),t_{b}=\hat{F}^{-1}_{j}(\tau_{b}|x)$ are the two closest values to $t$ such that $t_{a}\leq t\leq t_{b}$. Then, $\hat{F}_{j}(t|x)=\tau_{a}+\frac{\tau_{b}-\tau_{a}}{t_{b}-t_{a}}(t-t_{a})$.
@@ A.3 Estimation of the covariate-adjustment term
I fit all models in R. For quantile neural networks, I use the package qrnn with 3 hidden nodes, in 1 trial and a maximum of 100 iterations. Other parameters are the default choice. For quantile random forests, I use the package ranger with 1,000 trees and otherwise default hyperparameters. For all the regression methods, I use the package mlr3 with default values of hyperparameters. The packages for all models, used through mlr3 are: random forests (ranger), extreme gradient boosting (xgboost), support vector machine (e1071), neural networks (nnet), elastic net (glmnet). The constant/none model takes $s(x)=0$.
For choosing the best model in the application of Section 6, I use cross-validation inside each fold. That is, for each fold, I fit all the models in a 10-fold cross-validation and choose the model that gives the largest (smallest) lower (upper) bound. Note that the test set is never used in this approach.
@@
A.4 Definitions Ommited from the Main Text
@@
A.5 Proofs of the Main Results
@@
$\hskip 20.0pt-P\left(\hat{F}_{0,L}(\hat{t}_{max})-F_{0,L}(\hat{t}_{max})>c_{\alpha,0}\right)$ (A.1)
@@
$\hskip 15.0pt-\mathbb{E}\left[P\left(\sup_{t}\left(\hat{F}_{0,L}(t)-F_{0,L}(t)\right)>c_{\alpha,0}\Bigm{|}\mathcal{I}_{A}\right)\right]$ (A.2)
$\geq 1-\mathbb{E}\left[\exp{(-2|\mathcal{I}^{M}_{1}|c_{\alpha,1}^{2})}\right]-\mathbb{E}\left[\exp{(-2|\mathcal{I}^{M}_{0}|c_{\alpha,0}^{2})}\right]$ (A.3)
@@
(A.1) holds since $\hat{F}_{0,L}(t)-F_{0,L}(t)$ is symmetric around zero for all $t$. (A.2) holds by the law of iterated expectations, and note the expected values are in respect to $\mathcal{I}_{A}$. (A.3) holds by the one-sided DKW inequality as in Massart (1990). Note the use of sample splitting and conditioning on the auxiliary set $\mathcal{I}_{A}$ is important since it ensures that $(Y_{i},\hat{s}_{L}(X_{i}),D_{i})_{i\in\mathcal{I}^{1}_{M}\cup\mathcal{I}^{0}_{M}}$ are iid, assumption required for the DKW inequality.
@@
Theorem 3 is restated as Theorem A.1:
Theorem A.1 (Asymptotic distribution of cross-fitting estimators).
@@
$=\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(\hat{t}_{A},s_{A,P_{n}})-F_{P_{n},j}(\hat{t}_{A},s_{A,P_{n}})\right)+o_{P_{n}}(1)$ (A.4)
$=\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(t_{A,P_{n}},s_{A,P_{n}})-F_{P_{n},j}(t_{A,P_{n}},s_{A,P_{n}})\right)+o_{P_{n}}(1)$ (A.5)